OpenStudy (anonymous):

solve the proportion 12/5x-3 = 6/11 i knw the answer is 5 but how does it get solved??

OpenStudy (rulnick):

12/(5x-3) = 6/11 x=5.

OpenStudy (rulnick):

cross mult: 12(11) = (5x-3)6

OpenStudy (rulnick):

132 = 30x-18 150 = 30x 5 = x done

OpenStudy (anonymous):

thanx =)